a.) Calculate the molarity of 0.150 mol of Na2S in 1.10 L of solution?

b.) calculate the molarity of 32.9g of MgS in 763 mL of solution?


1
Expert's answer
2021-02-22T06:09:30-0500

a) molarity = moles / volume

Molarity = 0.150/1.10 = 0.136M

b) moles = 32.9/56.38 = 0.583moles

Volume = 763ml = 0.763L

Molarity = moles / volume

Molarity = 0.583/0.763 = 0.764 M